# Characterizing the MrDeepFakes Sexual Deepfake Marketplace

### v2

@@ 5.2.3 Requests for Deepfakes Targeting Children
Hello friends. Today I need 1 picture. There will be 2 boys. These need to be placed in 18+ pictures with me you can quickly reach email address [redacted]
@@
While this request did not receive any replies, it was also not taken down or given a warning by a site moderator like many other posts on the site. It is unclear whether the request was fulfilled through private communication.
@@ 5.3 Creators and Sellers
Over 1,800 users have uploaded deepfake video content on MrDeepFakes, though we note that $657$ users (0.1%) uploaded over 95% of the site’s video content, inline in line with the distribution of user participation on other social media platforms [nielsen2006participation]. Around 228 users have responded to requests for free or paid deepfake videos, of which 55 are “verified” sellers (Section 5.3.2); however, this is likely an underestimate of the total number of users creating videos since a significant fraction of communication occurs over private channels. 94.7% of paid and 72.9% of free requests are replied to, and the median time times to receive a response are 16 hours and 5 days, respectively.
Sellers responded to an average eight posts, but this is heavily skewed with a handful (10–15) of sellers responding to an outsized number of requests. One seller responded to 531 requests alone. In these cases, sellers appear to aggressively respond to most posts with generic replies like “PM sent,” “Please check your PM, thank you,” and Here is a sample of my work: […] dm me on discord: […]”
@@ 5.3.1 Marketing and Differentiation
Sellers commonly showcase their abilities to potential buyers by posting their own deepfake videos on MrDeepFakes’ s tube site. In paid request threads, 365 (18.9%) of deepfake creatorssellers’ replies included links to deepfake videos on the tube site as well as embedded screenshots or thumbnails from past work. Because deepfake approaches are rapidly evolving and creators want to showcase only their best work, this also appears to lead to churn in hosted videos. We observed that $1948$ of $42986$ video URLs were inaccessible; 75.1% of deleted videos indicated that they had been deleted upon their creator ’s creator request. Anecdotally, we find indications of increasing specialization among deepfake creators, like we observed with creators dedicated to deepfaking Guatemalan newscasters (Section 4.3). In response to a request seeking “somebody who would specialize [in] classic (1970s and 1980s)” deepfakes, a verified creator, moderator, and staff member responded:
I’d love to, but my thing would probably be modern celebs in 70s porn…Feel free to suggest a list of celebs…and if I’m inspired by one, I may do it.
@@ 5.3.2 Verified Creators
Deepfake creators can apply to become Verified Video Creators, who receive a “trusted” badge indicating that the user creates their own deepfake content and are able to monetize on the platform itself (e.g., paid video downloads, cryptocurrency payouts, advertising privileges in forum threadsforum advertising privileges). Verified creators also bypass manual moderation of their content content moderation, which site administrators claim is a process that all uploads from unverified users undergo and is expected to take takes several days. Further, platform affordances incentivize verification: the tube site showcases verified creators’ videos, and a verified creator’s verification status appears as a colorful label alongside their forum posts. The application process aims to certify that they generate their own content. To demonstrate their required technical knowledge for deepfake creation, applicants must provide evidence that they possess datasets or images of applications they use; they must also describe their creation process. Beyond proving their capabilities, the applicant must also have reasonable standing within the MrDeepFakes community, as they are required to link to their tube site profile with at least three “active” videos (i.e., they have not been removed by moderators). Their account must also be e-mail verified, at least 30 days old, and cannot have received warnings from moderators.
@@ 5.4 Cost of a Custom Deepfake
The average price advertised by sellers is \$87.50; 35.1% of quoted prices are at least \$100, and 12.3% are at least \$200. The range in prices is significant: the most inexpensive least expensive advertised service was \$1 for a one-minute video with a 30-minute turnaround time with one condition: “No [blowjobs].” Conversely, the highest price discussed was \$$1500$. For 31 of the 58 request posts with explicit prices, the length of the video was included as a parameter; the length of a deepfake video is loosely correlated with the commission price ($R=0.29$). From the limited visibility of a public vantage point, we cannot meaningfully determine if prices have changed over time. However, we do find anecdotes from community leaders that describe how deepfake technology has shifted the economy of the broader and longer-standing “faking” market:
Manual faking might be dead, faking in general isn’t, AI has made it very easy to the point making money off them is hard (it always was but now [because of] stricter laws, you already can’t monetize celebrity fakes on just about every patreon/onlyfans type of site), on the other hand AI opens doors to certain new things, because you can make content faster, easier, people who are creative but lacked certain skills or time can now produce content and make money of it (you do not have to make fakes of celebs, you can do many things with deepfacelab, roop, stable diffusion and ebsynth).
@@ 5.4.1 Factors Affecting Price
Several buyers noted the factors that affected the price that they were willing to pay; they were willing to pay more for videos with more than one target depicted, if the angles of the video are were complicated, or if their target was male. Because women are predominantly targeted by sexual deepfakespredominantly women are targeted by deepfake NCII, facesets and information regarding sex worker “body matches” for men are comparatively rarer. Several buyers described the factors they considered when choosing a seller to fulfill their request. Most commonly (12% of sampled requests), buyers explicitly stated that they sought high-quality deepfakes where a deepfake’s quality is typically defined by its resolution or the degree of achieved realism. Other factors included the estimated production turnaround time (1.9%) and the “professionalism” of the seller (0.8%). To proactively vet sellers, 12 requests asked interested creators to link to a portfolio of their past work; two other requests specified they were looking for a verified video creator (Section 5.3.2) or one with good reviews on the dedicated forum thread.
@@
Beware of cheap offers too, anyone saying they’ll do it for 10–50\$ can’t possibly deliver a quality product. (not naming anyone but I know one of the guys that often posts offers here made such a crap job for one of my clients I’d honestly ban you for this shit).
@@
The overall market for deepfakes has thus split into two tiers: (1) low-quality and fast “cheapfakes” and (2) high-end, custom “art.” The emergence of commodity deepfake applications (e.g., those marketed for face swapping, undressing, nudifying, etc. capabilities) [westerlund2019deepfake, cbs2019cheapfake, paris2019cheapfake] caters to the former, while seller scrutiny of cheapfake offers situates the paid requests marketplace on MrDeepFakes as offering the latter. These higher-end figures highlight the technical challenges posed in deepfake creation, which appear to require the expertise of a highly-skilled professional or the money to pay one.
@@
The only solution is find some guy studying to work for a serious SFX company who does TV and movies, and give him \$500 to \$1,500 I guess, but it’s probably even more than this…matching the light and integrat[ing] a body part so it feels like one with a real body is like 2 full weeks of work minimum, and for someone who knows what he’s doing.
@@
Another moderator agreed with this evaluation, advising the buyer to either “prepare plenty of bitcoins” or to choose a simpler scene, implying that such a difficult request would likely be prohibitively expensive and require niche expertise.
We find that this can occur even earlier between potential buyers and sellers than at sellers potentially establish private communication channels with potential buyers even before the price negotiation step. Of the $32284$ videos that have non-null descriptions, using regex and keyword search, we identify that 24.9% of the video descriptions provide viewers with contact information or otherwise solicit private communication. This positions videos uploaded to the tube site as advertisements themselves. Offers to communicate via first-party direct messages on the MrDeepFakes site itself are found in 8.9% of videos. Third-party platforms are also popular; the most common third-party platforms for direct messages include Telegram, Discord, and e-mail. We note limited use of third-party platforms that offer end-to-end encryption (E2EE) channels, identifying three references to Proton Mail addresses and 21 references to WhatsApp contact information. While Telegram offers E2EE in its “Secret Chat” mode, the default is not.
@@ 5.4.2 Additional Revenue Streams
Verified creators on MrDeepFakes can accept donations, sell video downloads (versus online playback), and offer technical support, and provide paid technical support or other resources through MrDeepFakes tokens, a first-party digital currency where the lowest denomination of 100 tokens can be bought using \$1.99 worth of BTC, or directly through their Bitcoin address. After inspecting the MrDeepFakes token purchase process, we found that token purchases occur only through a third-party cryptocurrency payment processor. While Bitcoin transactions are public, we would only know what address(es) to inspect if we purchased tokens ourselves, directly contributing to site revenue. Further, without making repeated token purchases over time, we would not be able to determine if transactions are split across different addresses. Therefore, we chose to not pursue further analysis of MrDeepFakes token purchase volumes. Though we cannot measure how many users pay to download video content, we observe that the advertised download price varied greatly, ranging from \$2 to \$32 in tokens.
Similarly, we collected and attempted analysis for 292 unique Bitcoin addresses provided in designated video metadata fields for direct donations. However, because of cryptocurrency transaction anonymity, we could not determine which transactions were related to deepfake services. Bounding transactions by associated MrDeepFakes accounts’ creation timestamps, we did find that 20% of addresses received at least \$1,000. Ultimately, we note that these estimates are too loosely bounded to make verified claims about deepfake-related revenue for these addresses. In addition, video descriptions on the tube site commonly ($15039$ of $42986$ videos, 35%) included third-party links. About half (49%) are to payment sites or subscription signups on both mainstream (e.g., PayPal, Ko-Fi, Patreon) and adult content (e.g., OnlyFans, FansMine, Fan-Topia) platforms.
@@ 6 Deepfake Video Creation
In addition to serving deepfake video content and providing a marketplace for buyers and sellers, MrDeepFakes also hosts discussions about the creation of deepfake videos. Prior work has established that while nonexperts can successfully create deepfakes, the ability to create realistic deepfakes (e.g., not suspicious to a human or software detector) requires deep expertise using complex, difficult-to-use software [mehta2023can]. An active community has emerged on MrDeepFakes to facilitate this process, using the forum to discuss best practices, offer tips and suggestions on improving deepfake quality, and provide detailed guides. In this section, we analyze forum discussion and the most popular guides for deepfake video production to document the requirements, challenges, and community solutions for the deepfake creation pipeline.
We specifically analyzed
@@ 6.1 Methodology
We specifically perform content analysis the “DeepFake Video Creation Tools” section of the forum, which is designated for members to discuss how to create deepfakes and also hosts guides and tutorials for deepfake creation, authored by expert community members, forum staff, administrators, moderators, and verified creators. The section contained 56.9% of all threads and 58.8% of all posts on the platform. We began by manually examining the most active guide threads (by number of replies), as well as guides linked in a pinned “quick start” post for newcomers. From there, we analyzed other linked threads, excluding topics unrelated to sexual deepfake videos, which uncovered 25 guide and tutorial threads containing $1229$ posts. We annotated discussions for troubleshooting topics or resource requirements for deepfake creation, surfacing three major themes: (1) access to facial image data, (2) software proficiency, and (3) powerful hardware. We discuss the challenges posed by each and how the MrDeepFakes community has developed solutions for each of them.
While we attempted large-scale, automated analysis of forum posts to understand the deepfake video creation process through BERTopic topic modeling, we encountered several obstacles. First, the topics that emerged did not provide insight beyond what the structure of the forum already offered (e.g., clusters about facesets, tools for deepfake video creation). Second, this approach fails to scale the process of revealing interesting thematic nuance within clusters, requiring manual annotation of topic clusters to effectively apply this method. Therefore, we ultimately pursued our manual approach.
@@ 6.2 Facial Image Data
Crowdsourced Facesets.
Crowdsourced Facesets.When facial image data is limited, users turn to the forum to request and share “facesets,” which are curated dumps of facial image data for a particular target. We find that facesets are often freely shared, as only two of the 711 threads that pertain to sharing facesets were labeled as paid solicitations. Of these threads, 81.6% contained links to third-party file sharing sites (Appendix B). In comparison, 326 threads request the creation of new facesets to target an individual; however, only 39.6% of these threads are replied to and only 8% are returned a link to a file sharing domain. Only three new faceset requests were paid offers. Experts’ discussion and purchase offers of facesets suggest that the curation of high quality facesets is labor-intensive and valuable:
I’m looking for someone able to make quality facesets…I’ll give you 3 to 4 to do every month. Regarding money, it will be paid in crypto…and depending on where you live it should be enough to pay some to all of your rent
@@
Anecdotally, we observe that buyers can purchase deepfakes at a lower cost if they are willing to provide their own faceset. We suspect that the lower success rate of finding willing community members to create new facesets for free is because of the labor-intensive process it entails. In short, not only do community members crowdsource in the face of obstacles posed by limited facial image data, but many community members are willing to share existing data for free.
@@ 6.3 Software
The FS dev team will absolutely not acknowledge any NSFW. The DFL dev has no issues with NSFW.
@@
DFL is currently very far ahead in dealing with face obstructions. FS is working on face obstructions, but nothing has been released yet.
@@
Both tools are free and open source but are complex to use. As a result, significant community effort is dedicated to streamlining and troubleshooting their usage, which we anecdotally find affects the adoption and popularity of deepfake software tools; troubleshooting comprised $3137$ (50.1%) of threads in the sexual deepfake video-related sections, representing 3,945 posts (41.6%) on the forum in aggregate. DFL is the subject of three pinned guides; the main DFL guide has the most views and replies (2M and 855, respectively), and the DFL repository creator and maintainer also appears to be an active member of the MrDeepFakes community, with an account adorned with a “Developer” badge. Several other authors of the DFL preprint [perov2021dfl] appear to be active members, with roles including developer, staff, administrator, and verified creator.
@@
To further investigate the role of open source software (OSS) tools, we extracted 936 Github links from forum discussion, uncovering 141 repositories. For each repository, we manually identified the type of software (e.g., image/video creation, audio/voice tools, generic machine learning libraries, system monitoring tools, datasets, etc.), whether the repository was a fork of another repository, whether the tool is permissive of NSFW content, and if the repository has any associated academic papers. 35 repositories (25%) are forks, of which 13 are forks of DFL, and 9 are forks of Roop, a “one-click face swap” library. Beyond just Roop, 17 (12%) repositories are a new class of tools that we term “limited-data” tools, which are designed to produce deepfakes without substantial data. The class has emerged in response to the obstacles posed by collecting sufficient quantities of facial image data, with some claiming to be able to do so with a single facial image.
Roop, Rope, and SwapFace.
@@
Roop, Rope, and SwapFace.Two “limited-data” tools stand out on MrDeepFakes: Roop/Rope, which have a moderator-written guide linked in the quick start, and SwapFace, which is the subject of the eighth-most active thread on the forum. In terms of output fidelity, some members praise the tools’ potential to produce high quality deepfakes for simple videos, while others lament their inability to do so for complicated videos. This parallels the striation of the deepfake market into demands for two categories: a lower-tier for “cheapfakes” and a higher-tier for complex scenes that require deeper expertise and extensive image data. The former is handled by these limited-data tools, whose output quality suffers for more complex tasks; buyers of the latter must then turn to the marketplace to pay an expert for their services.
@@
Roop is a deprecated open-source project that contains controls to censor outputs that are potentially NSFW. Rope forks Roop, removing safety filters and enhancing its usability for deepfake creation with a GUI. Of the nine Roop forks, seven (one of which is Rope) circumvent Roop’s SFW protections. Additionally, we find that MrDeepFakes members engineered NSFW workarounds to three of the nine repositories that self-identify as SFW open source limited-data tools (i.e., they have safeguards in place for NSFW outputs). Even with the best intentions from OSS maintainers of “SFW deepfake generators,” our results suggest these are being easily bypassed by expert deepfake creators, who are willing to share their solutions with the larger community.
Research Tools.
@@
Research Tools.In our annotation of 141 linked Github repositories, we identified that 60 repositories (including forked repositories) were directly linked to 43 academic papers; in aggregate, these open source research code repositories were referenced in the forum over 400 times. The associated papers range in publication date: the oldest is a 2001 paper that achieves color correction between source and destination images using statistical analysis [reinhard2001image], and the six most recent papers were published in 2023, the year we collected our forum data [liu2023dfl, radford2023robust, bain2023whisperx, an2023pano, pan2023drag, yuan2023reliableswap]. Of these papers, CVPR is by far the most prevalent venue, accounting for 13 of the 43 papers (30.2%); the remaining 30 papers are spread across 21 other venues, including well-known conferences ICCV, ECCV, NeurIPS, ICML, and SIGGRAPH. The frequent and up-to-date references to reputable academic publications demonstrates that the MrDeepFakes community follows cutting-edge research.
@@
The For the community favorite, DeepFaceLab DFL (the Github repository was linked 203 times), the corresponding paper was initially published as an ArXiv preprint in 2020 [perov2021dfl] and later published in the peer-reviewed journal Pattern Recognition in 2023. The delta between the two versions largely focuses on potential abuses of DeepFaceLabDFL, which both versions claim is most commonly used by VFX artists. In addressing the ethical issues of advancing deepfake technology, the authors “admit that deepfake techniques may affect public discourse quality and infringe upon the citizens’ portrait rights” when weaponized for “misinformation, manipulation, harassment, [or] persuasion.” Despite this, the authors argue that suppressing the publication would not stop its development and believe that it was their “responsibility to publish DeepFaceLab to the academia community formally.”” [liu2023dfl]. Neither version directly addresses the use or abuse of deepfake technology to generate intimate sexual imagery. In August 2024, the repository for DeepFaceLab DFL was disabled by Github for ToS GitHub for terms of service (ToS) violations.

### v3

@@ 5.2.3 Requests for Deepfakes Targeting Children
While this thread did not receive any replies, it was also not taken down or given a warning by a site moderator, like many other posts on the site. It is unclear whether the request was fulfilled through private communication.
@@ 5.3 Creators and Sellers
Over 1,800 users have uploaded deepfake video content on MrDeepFakes. We note that $657$ users (0.1%) uploaded over 95% of the site’s video content, in line with the distribution of user participation on other social media platforms [42]. Around 228 users have responded to requests for free or paid deepfake videos, of which 55 are “verified” sellers (Section 5.3.2); however, this likely underestimates the total number of users creating videos, since a significant fraction of communication occurs over private channels. 94.7% of paid and 72.9% of free requests are replied to, and the median times to receive a response are 16 hours and 5 days, respectively.
Sellers responded to an average eight posts, but this is heavily skewed with a handful (10–15) of sellers responding to an outsized number of requests. One seller responded to 531 requests alone. In these cases, sellers appear to aggressively respond to most posts with generic replies like “PM sent,” “Please check your PM, thank you,” and “Here is a sample of my work: […] dm me on discord: […]”
@@ 5.3.1 Marketing and Differentiation
Sellers commonly showcase their abilities to potential buyers by sharing their own deepfake videos on MrDeepFakes’ tube site. In paid request threads, 365 (18.9%) of deepfake sellers’ replies included links to deepfake videos on the tube site as well as embedded screenshots or thumbnails from past work. Because deepfake approaches are rapidly evolving and creators want to showcase their best work, this also appears to lead to churn in hosted videos. We observed that $1948$ of $42986$ video URLs were inaccessible; 75.1% of deleted videos indicated that they had been deleted upon creator request. Anecdotally, we find increasing specialization among deepfake creators, like we observed with creators dedicated to deepfaking Guatemalan newscasters (Section 4.3). In response to a request seeking “somebody who would specialize [in] classic (1970s and 1980s)” deepfakes, a verified creator, moderator, and staff member responded:
@@ 5.3.2 Verified Creators
Deepfake creators can apply to become Verified Video Creators, who receive a “trusted” badge indicating that the user creates their own deepfake content and is able to monetize on the platform (e.g., paid video downloads, cryptocurrency payouts, forum advertising privileges). Verified creators also bypass content moderation, which site administrators claim is a process that all unverified users’ uploads undergo and takes several days. Further, platform affordances incentivize verification: the tube site showcases verified creators’ videos, and a verified creator’s verification status appears as a colorful label alongside their forum posts. The application process aims to certify that they generate their own content. To demonstrate their required technical knowledge for deepfake creation, applicants must provide evidence that they possess datasets or images of applications they use, and they must also describe their creation process. Beyond proving their capabilities, the applicant must also have reasonable standing within the MrDeepFakes community, as they are required to link to their tube site profile with at least three “active” videos (i.e., not removed by moderators). Their account must also be e-mail verified, at least 30 days old, and cannot have received warnings from moderators.
@@ 5.4 Cost of a Custom Deepfake
The average price advertised by sellers is \$87.50; 35.1% of quoted prices are at least \$100, and 12.3% are at least \$200. The range in prices is significant: the least expensive advertised service was \$1 for a one-minute video with a 30-minute turnaround time with one condition: “No [blowjobs].” Conversely, the highest price discussed was \$$1500$. For 31 of the 58 request posts with explicit prices, the length of the video was included as a parameter; the length of a deepfake video is loosely correlated with the commission price ($R=0.29$). From the limited visibility of a public vantage point, we cannot meaningfully determine if prices have changed over time. However, we do find anecdotes from community leaders that describe how deepfake technology has shifted the economy of the broader and longer-standing “faking” market:
@@ 5.4.1 Factors Affecting Price
Several buyers noted the factors that affected the price that they were willing to pay; they were willing to pay more for videos with more than one target depicted, if the angles of the video were complicated, or if their target was male. Because predominantly women are targeted by deepfake NCII, facesets and information regarding sex worker “body matches” for men are comparatively rarer. Several buyers described the factors they considered when choosing a seller to fulfill their request. Most commonly (12% of sampled requests), buyers explicitly stated that they sought high-quality deepfakes where a deepfake’s quality is typically defined by its resolution or the degree of achieved realism. Other factors included the estimated production turnaround time (1.9%) and the “professionalism” of the seller (0.8%). To proactively vet sellers, 12 requests asked interested creators to link to a portfolio of their past work; two other requests specified they were looking for a verified video creator (Section 5.3.2) or one with good reviews on the dedicated forum thread.
@@
The overall market for deepfakes has thus split into two tiers: (1) low-quality and fast “cheapfakes” and (2) high-end, custom “art.” The emergence of commodity deepfake applications (e.g., those marketed for face swapping, undressing, nudifying, etc. capabilities) [66, 48, 47] caters to the former, while seller scrutiny of cheapfakes situates the paid requests marketplace on MrDeepFakes as offering the latter. These higher-end figures highlight the technical challenges posed in deepfake creation, which appear to require the expertise of a highly-skilled professional or the money to pay one.
@@
Another moderator agreed with this evaluation, advising the buyer to either “prepare plenty of [B]itcoins” or choose a simpler scene, implying that such a difficult request would likely be prohibitively expensive and require niche expertise.
We find that sellers potentially establish private communication channels with potential buyers even before the price negotiation step. Of the $32284$ videos that have non-null descriptions, using regex and keyword search, we identify that 24.9% of the video descriptions provide viewers with contact information or otherwise solicit private communication. This positions videos uploaded to the tube site as advertisements themselves. Offers to communicate via first-party direct messages on the MrDeepFakes site are found in 8.9% of videos. Third-party platforms are also popular; the most common third-party platforms for direct messages include Telegram, Discord, and e-mail. We note limited use of third-party platforms that offer end-to-end encryption (E2EE) channels, identifying three references to Proton Mail addresses and 21 references to WhatsApp contact information. While Telegram offers E2EE in its “Secret Chat” mode, the default is not.
@@ 5.4.2 Additional Revenue Streams
Verified creators on MrDeepFakes can accept donations, sell video downloads (versus online playback), and offer paid technical support or other resources through MrDeepFakes tokens, a first-party digital currency where the lowest denomination of 100 tokens can be bought using \$1.99 worth of BTC, or directly through their Bitcoin address. After inspecting the MrDeepFakes token purchase process, we found that token purchases occur only through a third-party cryptocurrency payment processor. While Bitcoin transactions are public, we would only know what address(es) to inspect if we purchased tokens ourselves, directly contributing to site revenue. Further, without making repeated token purchases over time, we would not be able to determine if transactions are split across different addresses. Therefore, we chose to not pursue further analysis of MrDeepFakes token purchase volumes. Though we cannot measure how many users pay to download video content, we observe that the advertised download price varied greatly, ranging from \$2 to \$32 in tokens.
Similarly, we collected and attempted analysis for 292 unique Bitcoin addresses provided in designated video metadata fields for direct donations. However, because of cryptocurrency transaction anonymity, we could not determine which transactions were related to deepfake services. Bounding transactions by associated MrDeepFakes accounts’ creation timestamps, we did find that 20% of addresses received at least \$$1000$. Ultimately, we note that these estimates are too loosely bounded to make verified claims about deepfake-related revenue for these addresses. In addition, video descriptions on the tube site commonly ($15039$ of $42986$ videos, 35%) included third-party links. About half (49%) are to payment sites or subscription signups on both mainstream (e.g., PayPal, Ko-Fi, Patreon) and adult content (e.g., OnlyFans, FansMine, Fan-Topia) platforms.
@@ 6 Deepfake Video Creation
In addition to serving deepfake video content and providing a marketplace for buyers and sellers, MrDeepFakes also hosts discussions about the creation of deepfake videos. Prior work has established that while nonexperts can successfully create deepfakes, the ability to create realistic deepfakes (e.g., not suspicious to a human or software detector) requires deep expertise using complex, difficult-to-use software [39]. An active community has emerged on MrDeepFakes to facilitate this process, using the forum to discuss best practices, offer tips and suggestions on improving deepfake quality, and provide detailed guides. In this section, we analyze forum discussion and the most popular guides for deepfake video production to document the requirements, challenges, and community solutions for the deepfake creation pipeline.
@@ 6.1 Methodology
We specifically perform content analysis on the “DeepFake Video Creation Tools” section of the forum, which is designated for members to discuss how to create deepfakes and also hosts guides and tutorials for deepfake creation, authored by expert community members, forum staff, administrators, moderators, and verified creators. The section contained 56.9% of all threads and 58.8% of all posts on the platform. We began by manually examining the most active guide threads (by number of replies), as well as guides linked in a pinned “quick start” post for newcomers. We then analyzed other linked threads, excluding topics unrelated to sexual deepfake videos, which uncovered 25 guide and tutorial threads containing $1229$ posts. We annotated discussions for troubleshooting topics or resource requirements for deepfake creation, surfacing three major themes: (1) access to facial image data, (2) software proficiency, and (3) powerful hardware. For each, we discuss the challenges posed and how the MrDeepFakes community has developed corresponding solutions.
We attempted large-scale, automated analysis of forum posts to understand the deepfake video creation process through BERTopic topic modeling. However, we encountered several obstacles. First, the generated topic clusters did not provide insight beyond the existing structure of the forum sections (e.g., posts discussing facesets, tools for deepfake creation, etc.). Second, this approach fails to scale revealing thematic nuances within clusters, requiring manual annotation of topic clusters to effectively apply this method. Therefore, we ultimately pursued our manual approach.
@@ 6.2 Facial Image Data
Crowdsourced Facesets. When facial image data is limited, users turn to the forum to request and share “facesets,” which are curated dumps of facial image data for a particular target. We find that facesets are often freely shared, as only two of the 711 threads that pertain to sharing facesets were labeled as paid solicitations. Of these threads, 81.6% contained links to third-party file sharing sites (Appendix B). In comparison, 326 threads request the creation of new facesets to target an individual; however, only 39.6% of these threads are replied to, and only 8% are returned a link to a file sharing domain. Only three new faceset requests were paid offers. Experts’ discussion and purchase offers of facesets suggest that the curation of high-quality facesets is labor-intensive and valuable:
@@
Anecdotally, we observe that buyers can purchase deepfakes at a lower cost if they are willing to provide their own faceset. We suspect that the lower success rate of finding willing community members to create new facesets for free is because of the requisite labor. In short, not only do community members crowdsource in the face of obstacles posed by limited facial image data, but many community members are willing to share existing data for free.
@@ 6.3 Software
Both tools are free and open source but are complex to use. As a result, significant community effort is dedicated to streamlining and troubleshooting their usage, which we anecdotally find affects the adoption and popularity of deepfake software tools. Troubleshooting comprised $3137$ (50.1%) of threads in the sexual deepfake video-related sections, representing 3,945 posts (41.6%) on the forum in aggregate. DFL is the subject of three pinned guides; the main DFL guide has the most views and replies (2M and 855, respectively), and the DFL repository creator and maintainer also appears to be an active member of the MrDeepFakes community, with an account adorned with a “Developer” badge. Several other authors of the DFL preprint [49] appear to be active members, with roles including developer, staff, administrator, and verified creator.
@@
To further investigate the role of open source software (OSS) tools, we extracted 936 Github links from forum discussion, uncovering 141 repositories. For each repository, we manually identified the type of software (e.g., image/video creation, audio/voice tools, generic machine learning libraries, system monitoring tools, datasets, etc.), whether the repository was a fork of another repository, whether the tool is permissive of NSFW content, and if the repository has any associated academic papers. 35 repositories (25%) are forks, of which 13 are forks of DFL, and nine are forks of Roop, a “one-click face swap” library. Beyond just Roop, 17 (12%) repositories are a new class of tools that we term “limited-data” tools, which are designed to produce deepfakes without substantial data. The class has emerged in response to the obstacles posed by collecting sufficient quantities of facial image data, with some claiming to be able to generate deepfakes with a single facial image.
@@
Roop, Rope, and FaceSwap. Two “limited-data” tools stand out on MrDeepFakes: Roop/Rope, which have a moderator-written guide linked in the quick start, and FaceSwap, which is the subject of the eighth-most active thread on the forum. In terms of output fidelity, some members praise the tools’ potential to produce high-quality deepfakes for simple videos, while others lament their inability to do so for complicated videos. This parallels the striation of the deepfake market into demands for two categories: a lower-tier for “cheapfakes” and a higher-tier for complex scenes that require deeper expertise and extensive image data. The former is handled by these limited-data tools, whose output quality suffers for more complex tasks; buyers of the latter must then turn to the marketplace to pay an expert for their services.
@@
Roop is a deprecated OSS project that contains controls to censor potentially NSFW outputs. Rope forks Roop, removing safety filters and enhancing usability for deepfake creation with a GUI. Of the nine Roop forks, seven (including Rope) circumvent Roop’s protections. Additionally, we find that MrDeepFakes members engineered NSFW workarounds to three of the nine repositories that self-identify as SFW (i.e., they have safeguards for NSFW outputs) limited-data OSS tools. Even with the best intentions from OSS maintainers of “SFW deepfake generators,” our results suggest these are easily bypassed by expert deepfake creators, who are willing to share their solutions with the larger community.
@@
Research Tools. In our annotation of 141 linked Github repositories, we identified that 60 repositories (including forks) were directly linked to 43 academic papers; in aggregate, these open source research code repositories were referenced in the forum over 400 times. The associated papers range in publication date: the oldest is a 2001 paper that achieves color correction between source and destination images using statistical analysis [56], and the six most recent papers were published in 2023, the year we collected our forum data [33, 54, 4, 2, 46, 69]. Of these papers, CVPR is by far the most prevalent venue, accounting for 13 of the 43 papers (30.2%); the remaining 30 papers are spread across 21 other venues, including well-known conferences ICCV, ECCV, NeurIPS, ICML, and SIGGRAPH. The frequent and up-to-date references to reputable academic publications demonstrates that the MrDeepFakes community follows cutting-edge research.
@@
For the community favorite, DFL (the Github repository was linked 203 times), the corresponding paper was initially published as an arXiv preprint in 2020 [49] and later published in the peer-reviewed journal Pattern Recognition in 2023. The delta between the two versions largely focuses on potential abuses of DFL, which both versions claim is most commonly used by “VFX artists.” In addressing the ethical issues of advancing deepfake technology, the authors “admit that deepfake techniques may affect public discourse quality and infringe upon the citizens’ portrait rights” when weaponized for “misinformation, manipulation, harassment, [or] persuasion.” Despite this, the authors argue that suppressing the publication would not stop its development and believe that it was their “responsibility to publish DeepFaceLab to the academia community formally” [33]. Neither version directly addresses the use or abuse of deepfake technology to generate sexual imagery. In August 2024, the repository for DFL was disabled by GitHub for terms of service (ToS) violations.